rectangle has a length of 12 what is the width if the perimeter is 60 units

Answer by tutoringisfun(17)

Good evening student. I hope that this solution is helpful to you.
One must use the following formula to answer the question:
p=2L+2w
as stated in the problem above p=60 and L=12 with this information we can now solve the problem.
1. substitute the known values into the perimeter formula for the variables
60=2(12)+2w
2. distribute the 2
60= 24+2w
3. subtract 24 from both sides of the equation:
60-24=24-24+2w
4. combine like terms on the left
36=2w

5. we want to isolate w, so we divide each side by 2

36/2=2w/2
the 2's on the right cancel each other out leaving:
36/2=w
6. solve for w
18=w
If you desire to check the answer, all you do is substitute all the known values into the formula and determine whether or not, they add up to 60.

2(12)+2(18)= 24+36=60
60=60 so the answer is correct.
I hope that this was helpful to you, , should you desire paid assistance, you can email me at the address listed in my profile or contact me via the following websites: Ask a Expert - Visit my Virtual Office at Presto Experts or via www.instaedu.com, or via my facebook page.